OpenStudy (anonymous):

Find the line of symmetry for the parabola whose equation is y = 2x^2 - 4x + 1. x = -1 x = 1 x = 2

OpenStudy (anonymous):

\[x=-\frac{b}{2a}\]

OpenStudy (anonymous):

if i knew how to do that i would haha

OpenStudy (anonymous):

\[y=ax^2+bx+c\] \[y=2x^2-4x+1\] what are \(a\) and \(b\) ?

OpenStudy (anonymous):

a is 2 and b is 4

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\(a=2\) is correct \(b=4\) is incorrect

OpenStudy (anonymous):

hmm?

OpenStudy (anonymous):

\[\large y=\color{red}ax^2+\color{blue}bx+c\]\[\large y=\color{red}2x^2\color{blue}{-4}x+1\]

OpenStudy (anonymous):

tbh ima fail this class

OpenStudy (anonymous):

yeah i guess is it not clear that the minus sign comes with the number, so that \(a=2,b=-4\) ?

OpenStudy (anonymous):

okay i get that now

OpenStudy (anonymous):

then your last job is to compute \(-\frac{b}{2a}\) with \(a=2,b=-4\)

OpenStudy (anonymous):

allright so it will be \[x= \frac{ -4 }{ 22}\]

OpenStudy (anonymous):

\(2a\) means \(2\times 2\) not \(22\)

OpenStudy (anonymous):

haha so it would b 4

OpenStudy (anonymous):

also you forgot that \(b=-4\) and not \(4\) yes the denominator is 4

OpenStudy (anonymous):

\[x=-\frac{ 4 }{ 4 }\] i hate algebra

OpenStudy (anonymous):

\[\huge -\frac{\color{blue}b}{2\times\color{red}a}\]

OpenStudy (anonymous):

\[\huge -\frac{\color{blue}{-4}}{2\times\color{red}2}\]

OpenStudy (anonymous):

yeah i hate algebra too, but this is arithmetic actually

OpenStudy (anonymous):

now the only question is, since we know \(2\times 2=4\) what is \[\large -\frac{-4}{4}\]

OpenStudy (anonymous):

answer is 1

OpenStudy (anonymous):

yes